# Regional chemical potential analysis for material surfaces

### v1

@@ I Introduction
Let us consider dividing the system into $R$ and $R^{\prime}$ spatial regions. The electron transfer occurs from one region to another through the interface situated in-between. The regional DFT [4] has revealed that even if the system comes to an equilibrium state by electron transfer, the RCPs $\mu_{R}$ and $\mu_{R^{\prime}}$ are not necessarily equal to each other, nor to the chemical potential of the whole system. This RCP inequality principle has been investigated for the electron transfer in a HeH+ molecule based on the regional DFT [4]. The local version of the RCP under a linear response approximation is proposed in Refs. [5, 6] in the context of the bond order analysis. The spatial distribution analysis of the local version of the RCP is demonstrated for a Pt cluster surface to find reactivity regions for hydrogen adsorption [7].
In this paper, we propose a more general RCP analysis method using the electronic state around Fermi surface. By applying the analysis method to material surfaces, it also reveals that even in the case of covalent bonding, the RCP analysis at the material surface is possible to discuss quantitatively how selectively atoms and molecules can be adsorbed on a surface. By replacing the adsorbed molecule with a probe commonly used in non-contact or near-contact Atomic Force Microscopy (AFM), the ease of molecular adsorption can be interpreted in terms of the strength of the interaction force between the probe tip atom and the material surface. When both the tip and surface are ionized, ionic interactions can be observed, whereas in the case of a neutral probe, such as a Si tip, chemical bonding forces can be detected [8]. In other words, in terms of AFM experiments, RCP has the potential to evaluate bonding forces between the AFM tip and the material surface.
This paper is organized as follows. In Secs. II.1 and II.2, the definitions of local RCP and a material surface are introduced in the framework of regional DFT and relativistic gauge invariant theory of quantum electrodynamics (QED). In Secs. II.3 and S1.3, the relation between local RCP and the local picture of covalent bonding is revealed using the electronic stress tensor and the regional energy density. In Sec. III, local RCP of a hydrogen atom model is evaluated to grasp a simple interpretation of the local RCP at the material surface. In addition, direct relation between the chemical bonding force and local RCP is revealed using H2 molecular models. After introducing an approximation method for local RCP of general materials based on DFT in Sec. IV, computational details and DFT results of various types of organic molecules and C(001) diamond surfaces are discussed in Secs. V and VI. The last section is devoted to the conclusions and perspectives.
@@ II.1 Regional chemical potential
where the total energy of the system $E=E_{R}+\displaystyle\sum_{R^{\prime}}E_{R^{\prime}(\neq R)}$, and the total number of electron $N=N_{R}+\displaystyle\sum_{R^{\prime}}N_{R^{\prime}(\neq R)}$. The presence of the second term, which represents the quantum interferece effect, prohibits the equality in between $\mu$ and $\mu_{R}$ [2]. Suppose that the interferece term can be ignored, it satisfies Sanderson’s principle of electronegativity equalization [10], $\mu=\mu_{R}=\mu_{R^{\prime}}=\cdots$.
To understand the physical meaning of $\mu_{R}$, let us consider the three-step processes of electron emission shown in Fig. 1 (a). (i) Electrons are emitted from the surface of the region $R_{i}$ to the hypothetical region outside $R$. (ii) After electrons are emitted from the surface, the potential generated from electrons in the region $R_{i}$ changes. (iii) The above change in potential causes electrons in region $R_{j\neq i}$ to flow into region $R_{i}$ until $\mu$ in region $R$ becomes constant. From Eq. (4), the RCP includes the excitation effects of electrons associated with the processes (i) and (ii), but note that the process (iii), which corresponds to electron redistribution, is not included.
@@
Since the RCP in Eq. (4) cannot be calculated exactly in DFT unless the region $R$ is isolated from the other region, we introduce the local version of the RCP as
$\mu_{R}(\vec{r})=\frac{\Delta\varepsilon(\vec{r})}{\Delta n(\vec{r})},$ (6)
where $\Delta\varepsilon(\vec{r})$ and $\Delta n(\vec{r})$ are variations of energy density and electron number density, respectively, under a given perturbative electron excitation. More detailed approximation expression will be given in Sec. IV. Similar to the ordinary chemical potential, which is known to describe the electron transferability, the local version of the RCP at the surface represents how much electrons are emitted or absorbed from the material surface. That will be discussed in Sec. S1.3. Before starting the discussion on $\mu_{R}(\vec{r})$ itself, we introduce the definitions of a material surface and the energy density, which are significant to evaluate $\mu_{R}(\vec{r})$ on the surface.
@@ II.2 Definition of a material surface
where $\hat{\psi}$ is the 4-component Dirac field operator. The gauge covariant derivative is defined as $\hat{D}_{ek}(x)=\partial_{k}+i\frac{Z_{e}e}{\hbar c}\hat{A}_{k}(x)$, where $Z_{e}=-1$, $e\,(>0)$ is the elementary charge, $m_{e}$ is the electron mass, $c$ is the speed of light in vacuum, and $\hat{A}_{k}(x)$ is the vector potential of the photon field operator. As we treat non-relativistic systems, we approximated the expressions of the first equation in the framework of the primary Rigged QED (PRQED) approximation [9]. Under the approximation, the 4-component $\hat{\psi}$ is decomposed into the large 2-component $\hat{\psi}^{L}$ and the small 2-component $\hat{\psi}^{S}\approx\frac{1}{2m_{e}c}i\hbar\sigma^{k}\hat{D}_{ek}\hat{\psi}^{L}$, where $\sigma^{k}$ are the $2\times 2$ Pauli matrices, and the spin-dependent terms are ignored after the 2-component decomposition.
Although the kinetic energy is always positive in classical mechanics, the value of $\langle\hat{T}_{e}\rangle$ in quantum theory is not always positive. The meaning of the sign of $\langle\hat{T}_{e}\rangle$ can be explained as follows. For example, when we assume that electrons are trapped inside a potential well, $\langle\hat{T}_{e}\rangle>0$ in the potential well while $\langle\hat{T}_{e}\rangle<0$ outside the well. While the region where $\langle\hat{T}_{e}\rangle>0$ can be understood as the electron is allowed to move from the classical mechanical view, the region where $\langle\hat{T}_{e}\rangle<0$ can be understood as a tunneling region from the quantum theoretical view. Therefore, the boundary $\langle\hat{T}_{e}\rangle=0$ represents a surface of the material. This definition of a surface of material is useful, since the zero kinetic energy density is determined by the solution of the electronic state without introducing any artificial and arbitrary parameters. The zero kinetic energy density isosurface is called “electronic interface” [9, 11]. For atoms and molecules, there are some electronic interfaces inside of inner electronic structures, since they have a shell structure. However, it is easy to imagine that the outermost electronic interface defines the shape of atoms and molecules. The shape of the electronic interface of atoms is investigated in detail in Ref. [11].
@@ II.3 Energy density and electronic stress tensor
The energy density $\varepsilon(\vec{r})$ is the 0-0 component of the symmetric energy momentum tensor density $T^{\mu\nu}(\vec{r})$, which is defined in the general relativistic QED theory [12, 13, 14, 15, 16, 17]. Suppose that the system is in an equilibrium stationary state in the Minkovsky space limit, virial theorem can be applied. Furthermore, under the Born-Oppenheimer approximation, the total energy of the Rigged QED system is given as [18, 17]
@@
$\approx\int\langle m_{e}c^{2}\hat{\psi}^{\dagger}\gamma^{0}\hat{\psi}\rangle d^{3}\vec{r},$ (10)
where the $\gamma^{0}$ is the $4\times 4$ gamma matrix. Under the PRQED approximation for taking the non-relativistic limit, the total energy minus the mass energy is reduced to
@@
The detailed derivation is given in Sec. S1B in the Supplemental Information(SI). The symmetric electronic stress tensor operator in the PRQED is defined as
@@
$\approx\frac{\hbar^{2}}{4m}\left[\hat{\psi}^{\dagger}(x)\hat{D}_{ek}(x)\hat{D}_{el}(x)\hat{\psi}(x)-\left(\hat{D}_{ek}(x)\hat{\psi}(x)\right)^{\dagger}\hat{D}_{el}(x)\hat{\psi}(x)\right]+h.c.,$ (13)
where $\hat{T}_{e}^{kl}(x)$ is the electronic symmetric energy momentum tensor density operator111 The scalar expression $\hat{T}_{e}$ denotes the kinetic energy density operator, while the second-rank tensor $\hat{T}_{e}^{kl}$ denotes the electronic symmetric energy momentum tensor operator. , and $x=(ct,\vec{r})$ The expression of Eq. (11) reveals that only the integral of the trace of the electronic stress tensor is needed to evaluate the total energy under the above approximations. Using Eq. (11), we introduce another energy density operator expression as
@@
Here, the $\varepsilon_{\tau}(\vec{r})$ is called “regional energy density” [9, 5, 17], where $\varepsilon_{\tau}(\vec{r})\neq\varepsilon_{\rm PRQED}(\vec{r})$, but $\int\varepsilon_{\tau}(\vec{r})d^{3}\vec{r}=\int\varepsilon_{\rm PRQED}(\vec{r})d^{3}\vec{r}$. The $\varepsilon_{\rm PRQED}(\vec{r})$ is the non-relativistic version of the original energy density $\varepsilon(\vec{r})$, which appeared in Eq. (9). The advantage of this regional energy density is that the $\varepsilon_{\tau}(\vec{r})$ can be obtained from electronic density matrices for the considered system without additional calculations for the complicated interaction energy. Furthermore, the electronic stress tensor density appeared in the $\varepsilon_{\tau}(\vec{r})$ is known to be significant for a discussion of chemical bonding theory as mentioned in the next section.
@@ II.4 Covalent bonding
Let us consider a system such as near-contact AFM measurements [19, 20, 21, 8], in which the interaction between the material surface and the probe tip gives rise to forces such as ionic bonding, covalent bonding, van der Waals interactions, etc. For ionic systems, it is easy to understand that the difference in the RCP between the material surface and the probe tip is a measure for the ionic bonding, since electrons tend to move from regions of higher chemical potential to regions of lower chemical potential as shown in Fig. 1 (b). What can be said about covalent bonding? Even if the ordinary chemical potentials are the same over the system, the RCPs can be different in each region. In such a case, it can be assumed that electrons in the highly electron-donating region, the higher $\mu_{R}(\vec{r})$ region, are more likely to be supplied to the bonding region and contribute to covalent bonding as shown in Fig. 1 (c).
Next, we connect this physical interpretation with a local mechanical picture of a covalent bonding [22]. Normally, the eigenvalues of the electron stress tensor density are negative because the electrons repel each other and try to diffuse. However, in a covalent bonding state, the electrons in the bonding region are in a tensile stress state; namely the inphase overlap of orbitals brings positive maximum eigenvalues of $\hat{\tau}_{e}^{Skl}(x)$. This is known as a local mechanical picture of covalent bonding [18] (see Fig. S1 in SI). In addition, the regional energy density $\varepsilon_{\tau}(\vec{r}\,)$, which is the trace of the stress tensor density $\tau_{e}^{Skl}(\vec{r}\,)$, increases around the bonding region in the tensile stress state. As can be seen from the calculations for a hydrogen molecule, covalent bonding increases the regional energy density in the bonding region and decreases the regional energy density near the nucleus (see Fig. S2 in SI). The total energy as an integral of the total space is then reduced, and the system is stabilized. This is a mechanism of the covalent bonding stabilization in terms of the regional energy density. Motivating this chemical bonding picture, we define the local RCP $\mu_{R}^{\tau}(\vec{r})$ and its inverse, “local density response function” $\xi_{R}^{\tau}(\vec{r})$, using the regional energy density as
@@
where $\xi^{\tau}_{R}(\vec{r}\,)$ represents how much electrons are transferred by the energy change. If $\xi^{\tau}_{R}(\vec{r}\,)$ is negative and has a large absolute value, it means that electrons are more likely to transfer from the region $R$ when the energy increases in the bonding region due to tensile stress during the covalent bond formation process. The stronger interaction makes the electrons transfer from the region $R$ easier. That increases the electron overlap, and it makes the interaction stronger. Although both $\mu^{\tau}_{R}(\vec{r}\,)$ and $\xi^{\tau}_{R}(\vec{r}\,)$ might be useful for analysis of electron transfer and chemical bonding, $\mu^{\tau}_{R}(\vec{r}\,)$ is adopted in this study.

### v2

@@ I Introduction
Let us consider dividing the system into $R$ and $R^{\prime}$ spatial regions. The electron transfer occurs from one region to another through the interface situated in-between. The regional DFT [4] has revealed that even if the system comes to an equilibrium state by electron transfer, the RCPs $\mu_{R}$ and $\mu_{R^{\prime}}$ are not necessarily equal to each other, nor to the chemical potential of the whole system. This RCP inequality principle has been investigated for the electron transfer in a HeH+ molecule based on the regional DFT [4]. The local version of the RCP under a linear response approximation is proposed in Refs. [5, 6] in the context of the bond order analysis. The spatial distribution analysis of the local version of the RCP is demonstrated for a Pt cluster surface to find reactivity regions for hydrogen adsorption [7], and for a carbon nanotube surface to investigate lithium atom adsorption [8].
In this paper, we propose a more general RCP analysis method using the electronic state around the Fermi surface. By applying the analysis method to material surfaces, we also reveal that even in the case of covalent bonding, the RCP analysis at the material surface allows us to discuss quantitatively how selectively atoms and molecules can be adsorbed on a surface. By replacing the adsorbed molecule with a probe commonly used in non-contact or near-contact Atomic Force Microscopy (AFM), the ease of molecular adsorption can be interpreted as the strength of the interaction force between the probe tip atom and the material surface. When both the tip and surface are ionized, ionic interactions can be observed, whereas in the case of a neutral probe, such as a Si tip, chemical bonding forces can be detected [9]. In other words, in terms of AFM experiments, RCP has the potential to evaluate bonding forces between the AFM tip and the material surface.
This paper is organized as follows. In Secs. II.1 and II.2, the definitions of local RCP and a material surface are introduced in the framework of regional DFT and relativistic gauge invariant theory of quantum electrodynamics (QED). In Secs. II.3 and II.4, the relation between local RCP and the local picture of covalent bonding is revealed using the electronic stress tensor and the regional energy density. In Sec. III, local RCP of a hydrogen atom model is evaluated to grasp a simple interpretation of the local RCP at the material surface. In addition, a direct relation between the chemical bonding force and the local RCP is revealed using H2 molecular models. After introducing an approximation method for local RCP of general materials based on DFT in Sec. IV, computational details and DFT results of various types of organic molecules and C(001) diamond surfaces are discussed in Secs. V and VI. The last section is devoted to the conclusions and perspectives.
@@ II.1 Regional chemical potential
where the total energy of the system $E=E_{R}+\displaystyle\sum_{R^{\prime}}E_{R^{\prime}(\neq R)}$, and the total number of electron $N=N_{R}+\displaystyle\sum_{R^{\prime}}N_{R^{\prime}(\neq R)}$. The presence of the second term, which represents the quantum interference effect, prohibits the equality in between $\mu$ and $\mu_{R}$ [2]. Suppose that the interference term can be ignored, it satisfies Sanderson’s principle of electronegativity equalization [11], $\mu=\mu_{R}=\mu_{R^{\prime}}=\cdots$.
To understand the physical meaning of $\mu_{R}$, let us consider the three-step processes of electron emission shown in Fig. 1 (a). Here, we consider the situation where an electron is emitted from a subdivided region $R_{i}$ , obtained by dividing the region $R$ into several parts. (i) Electrons are emitted from the surface of the region $R_{i}$ to the hypothetical region outside $R$. (ii) After electrons are emitted from the surface, the potential generated from electrons in the region $R_{i}$ changes. (iii) The above change in potential causes electrons in region $R_{j\neq i}$ to flow into region $R_{i}$ until $\mu$ in region $R$ becomes constant. From Eq. (4), the RCP includes the excitation effects of electrons associated with the processes (i) and (ii), but note that the process (iii), which corresponds to electron redistribution, is not included.
@@
Since the RCP in Eq. (4) cannot be calculated exactly in DFT unless the region $R$ is isolated from the other region, at first, we introduce the finite derivative expression of the RCP as
$\mu_{R}\approx\left(\frac{\Delta E_{R}}{\Delta N_{R}}\right)_{N_{R^{\prime}(\neq R)}}.$ (6)
When taking the limit in which the region $R$ becomes infinitesimally small, the local version of the RCP $\mu_{R}(\vec{r})$ at each point in space can be expressed as
$\mu_{R}(\vec{r})\approx\frac{\Delta\varepsilon(\vec{r})}{\Delta n(\vec{r})},$ (7)
where $\Delta\varepsilon(\vec{r})$ and $\Delta n(\vec{r})$ are variations of energy density and electron number density, respectively, under a given perturbative electron excitation.
More detailed approximation expression will be given in Sec. IV. Similar to the conventional chemical potential, which is known to describe the electron transferability, the local version of the RCP at the surface represents how much electrons are emitted from or adsorbed onto the material surface. That will be discussed in Sec. II.4. Before starting the discussion on $\mu_{R}(\vec{r})$ itself, we introduce the definitions of a material surface and the energy density, which are significant to evaluate $\mu_{R}(\vec{r})$ on the surface.
@@ II.2 Definition of a material surface
where $\hat{\psi}(x)$ is the 4-component Dirac field operator, $x=(ct,\vec{r})$, and $c$ is the speed of light in vacuum. The 4-component $\hat{\psi}(x)$ is decomposed into the large 2-component $\hat{\psi}^{L}(x)$ and the small 2-component $\hat{\psi}^{S}(x)$ in the Dirac representation as $\hat{\psi}(x)=\left(\begin{smallmatrix}\hat{\psi}^{L}(x)\\ \hat{\psi}^{S}(x)\end{smallmatrix}\right)$. The gauge covariant derivative is defined as $\hat{D}_{ek}(x)=\partial_{k}+i\frac{Z_{e}e}{\hbar c}\hat{A}_{k}(x)$, where $Z_{e}=-1$, $e\,(>0)$ is the elementary charge, $m_{e}$ is the electron mass, and $\hat{A}_{k}(x)$ is the vector potential operator. As we treat non-relativistic systems, we approximated the expressions of the first equation in the framework of the primary Rigged QED (PRQED) approximation [10]. Under the approximation, $\hat{\psi}^{S}(x)\approx\frac{1}{2m_{e}c}i\hbar\sigma^{k}\hat{D}_{ek}(x)\hat{\psi}^{L}(x)$, where $\sigma^{k}$ are the $2\times 2$ Pauli matrices, and the spin-dependent terms are ignored after the 2-component decomposition.
Although the kinetic energy is always positive in classical mechanics, the value of $\langle\hat{T}_{e}(x)\rangle$ in quantum theory is not always positive. The meaning of the sign of $\langle\hat{T}_{e}(x)\rangle$ can be explained as follows. For example, when we assume that electrons are trapped inside a potential well, $\langle\hat{T}_{e}(x)\rangle>0$ in the potential well, while $\langle\hat{T}_{e}(x)\rangle<0$ outside the well. While the region where $\langle\hat{T}_{e}(x)\rangle>0$ can be understood as the electron is allowed to move from the classical mechanical view, the region where $\langle\hat{T}_{e}(x)\rangle<0$ can be understood as a tunneling region from the quantum theoretical view. Therefore, the boundary $\langle\hat{T}_{e}(x)\rangle=0$ represents the surface of the material. This definition of the material surface is useful, since the zero kinetic energy density is determined by the solution of the electronic state without introducing any artificial or arbitrary parameters. The zero kinetic energy density isosurface is called “electronic interface” [10, 12]. For atoms and molecules, there are some electronic interfaces inside of inner electronic structures, since they have a shell structure. However, it is easy to imagine that the outermost electronic interface defines the shape of atoms and molecules. The shape of the electronic interface of atoms is investigated in detail in Ref. [12].
@@ II.3 Energy density and electronic stress tensor
The energy density $\varepsilon(\vec{r})$ is the 0-0 component of the symmetric energy momentum tensor density $T^{\mu\nu}(\vec{r})$, which is defined in the general relativistic QED theory [13, 14, 15, 16, 17, 18]. Suppose that the system is in an equilibrium stationary state in the Minkowski space limit. In that case, the virial theorem can be applied. Furthermore, under the Born-Oppenheimer approximation, the total energy of the Rigged QED system is given as [19, 18]
@@
$\approx\int\langle m_{e}c^{2}\hat{\psi}^{\dagger}(\vec{r})\gamma^{0}\hat{\psi}(\vec{r})\rangle d^{3}\vec{r},$ (11)
where $\gamma^{0}$ is the $4\times 4$ gamma matrix. Under the PRQED approximation for taking the non-relativistic limit, the total energy minus the mass energy is reduced to
@@
The detailed derivation is given in Sec. S2 in the Supplemental Information(SI). The symmetric electronic stress tensor operator in the PRQED is defined as
@@
$\approx\frac{\hbar^{2}}{4m_{e}}\left[\hat{\psi}^{L\dagger}\hat{D}_{ek}\hat{D}_{el}\hat{\psi}^{L}-\left(\hat{D}_{ek}\hat{\psi}^{L}\right)^{\dagger}\hat{D}_{el}\hat{\psi}^{L}\right]+h.c.,$ (14)
where $\hat{T}_{e}^{kl}(x)$ is the electronic symmetric energy momentum tensor density operator111 The scalar expression $\hat{T}_{e}$ denotes the kinetic energy density operator, while the second-rank tensor $\hat{T}_{e}^{kl}$ denotes the electronic symmetric energy momentum tensor operator. . The expression of Eq. (12) reveals that only the integral of the trace of the electronic stress tensor is needed to evaluate the total energy under the above approximations. Using Eq. (12), we introduce another energy density operator expression as
@@
Here, $\varepsilon_{\tau}(\vec{r})$ is referred to as the “regional energy density”[10, 5, 18]. Its definition differs from that of the non-relativistic version of the original energy density $\varepsilon(\vec{r})$ in Eq. (10), even when excluding the mass energy density. However, the energy obtained by integrating over the whole space excluding the mass energy is identical. The advantage of this regional energy density lies in the fact that $\varepsilon_{\tau}(\vec{r})$ can be derived directly from the electronic density matrices, without the need for additional intricate calculations of the interaction energies. Furthermore, the electronic stress tensor density appeared in $\varepsilon_{\tau}(\vec{r})$ is known to be significant for a discussion of chemical bonding theory as mentioned in the next section.
@@ II.4 Covalent bonding
Let us consider a system such as near-contact AFM measurements [20, 21, 22, 9], in which the interaction between the material surface and the probe tip gives rise to forces such as ionic bonding, covalent bonding, van der Waals interactions. For ionic systems, it is easy to understand that the difference in the RCP between the material surface and the probe tip is a measure for the ionic bonding, since electrons tend to move from regions of higher chemical potential to regions of lower chemical potential, as shown in Fig. 1 (b). What can be said about covalent bonding? Even if the ordinary chemical potentials are the same over the system, the RCPs can be different in each region. In such a case, it can be assumed that electrons in the highly electron-donating region, the higher $\mu_{R}(\vec{r})$ region, are more likely to be supplied to the bonding region and contribute to covalent bonding as shown in Fig. 1 (c).
Next, we relate this physical interpretation to a local mechanical picture of covalent bonding [23]. Normally, the eigenvalues of the electron stress tensor density are negative because the electrons repel each other and try to diffuse. However, in a covalent bonding state, the electrons in the bonding region are in a tensile stress state; namely the inphase overlap of orbitals brings positive maximum eigenvalues of $\langle\hat{\tau}_{e}^{Skl}(\vec{r})\rangle$. This is known as a local mechanical picture of covalent bonding [19] (see Fig. S1 in SI). In addition, the regional energy density $\varepsilon_{\tau}(\vec{r}\,)$, which is a half of the trace of the stress tensor density $\langle\hat{\tau}_{e}^{Skl}(\vec{r}\,)\rangle$, increases around the bonding region in the tensile stress state. As can be seen from the calculations for a hydrogen molecule, covalent bonding increases the regional energy density in the bonding region and decreases the regional energy density near the nuclei (see Fig. S2 in SI). The total energy as an integral over the whole space is then reduced, and the system is stabilized. This is a mechanism of the covalent bonding stabilization in terms of the regional energy density. Motivating this chemical bonding picture, we define the local RCP $\mu_{R}^{\tau}(\vec{r})$ and its inverse, “local density response function” $\xi_{R}^{\tau}(\vec{r})$, using the regional energy density as
@@
where $\xi^{\tau}_{R}(\vec{r}\,)$ represents how much electrons are transferred by the energy change. If $\xi^{\tau}_{R}(\vec{r}\,)$ is negative and has a large absolute value, it means that electrons are more likely to transfer from the region $R$ when the energy increases in the bonding region due to tensile stress during the covalent bond formation process. The stronger interaction facilitates the transfer of electrons from region $R$. That increases the electron overlap, and it makes the interaction stronger. Although both $\mu^{\tau}_{R}(\vec{r}\,)$ and $\xi^{\tau}_{R}(\vec{r}\,)$ might be useful for analysis of electron transfer and chemical bonding, $\mu^{\tau}_{R}(\vec{r}\,)$ is adopted in this study.